OpenStudy (supernova_sonntag):

So I need to understand how to solve this type of problem... y=(x+2)^2-6 y-2+x I think by...substitution? And I know you probably have to foil at some point but I have tried this problem so many times and I can't figure it out.

OpenStudy (anonymous):

do you mean y=2+x for the second function?

OpenStudy (anonymous):

And what are you trying to solve for?

OpenStudy (supernova_sonntag):

my bad I meant y-2=x silly shift key :) the directions only said solve...I am assuming it meant solve the system of equations

OpenStudy (xapproachesinfinity):

So we have \(y=(x+2)^2-6 \\ y-2=x\)

OpenStudy (xapproachesinfinity):

is that what yu have?

OpenStudy (anonymous):

I'm gonna let Mr. Mathlete take it from here

OpenStudy (supernova_sonntag):

Yes

OpenStudy (xapproachesinfinity):

let us substitute x for y-2 in equation #1

OpenStudy (xapproachesinfinity):

so to get \(y=(y-2+2)^2-6\)

OpenStudy (xapproachesinfinity):

you can see that i replaced x by y-2 right?

OpenStudy (supernova_sonntag):

Yeah

OpenStudy (xapproachesinfinity):

ok now just simplify that more \(y=y^2-6\)

OpenStudy (xapproachesinfinity):

-2 +2 cancel out

OpenStudy (xapproachesinfinity):

now it need more clean up \(y^2-y-6=0\)

OpenStudy (xapproachesinfinity):

i subtract y from both sides to get that

OpenStudy (xapproachesinfinity):

what kind of equation is that?

OpenStudy (supernova_sonntag):

Okay I see what you're doing, i'm not sure what you mean though, by what kind of equation?

OpenStudy (xapproachesinfinity):

i meant that's a quadratic equation, you should have faced it various times?

OpenStudy (supernova_sonntag):

Yes, but honestly I didn't recognize that; in my math class we don't really do much with the vocabulary for things, unfortunately.

OpenStudy (xapproachesinfinity):

vocabulary is really part of mathematics

OpenStudy (xapproachesinfinity):

we need a language to cross our ideas!

OpenStudy (supernova_sonntag):

Haha very true. So, is there anything more that needs to be done to the problem? Or, is that the type of answer they are looking for?

OpenStudy (xapproachesinfinity):

so back to our mean thing. \(y^2-y-6=0 \Longrightarrow (y-3)(y+2)=0 \)

OpenStudy (xapproachesinfinity):

that is called factoring and we are not done until we figure out what x is and y

OpenStudy (xapproachesinfinity):

now we pretty much close in finding y

OpenStudy (supernova_sonntag):

ooooh

OpenStudy (xapproachesinfinity):

any idea what to do with \((y-3)(y+2)=0\)

OpenStudy (supernova_sonntag):

Um to solve for y you set them both equal to zero?

OpenStudy (xapproachesinfinity):

yes

OpenStudy (supernova_sonntag):

okay so y=3 and y=-2 right?

OpenStudy (xapproachesinfinity):

yep!

OpenStudy (xapproachesinfinity):

so we got two values for y we take one value at a time and look for x

OpenStudy (xapproachesinfinity):

to find x we go back the equations we started with

OpenStudy (xapproachesinfinity):

what equation do you think is easier to spit out an answer for us

OpenStudy (supernova_sonntag):

the second one

OpenStudy (xapproachesinfinity):

yes

OpenStudy (xapproachesinfinity):

now pick one y value you found and plug it to that equation and find x

OpenStudy (supernova_sonntag):

so x=1 and -4?

OpenStudy (xapproachesinfinity):

oh you already tried both values of y

OpenStudy (xapproachesinfinity):

yes that good answers (1, 3) and (-4, -2)

OpenStudy (xapproachesinfinity):

the answer is two points

OpenStudy (supernova_sonntag):

Okay. Thank you so much, I understood this a while ago, but I completely forgot. You jogged my memory!!

OpenStudy (xapproachesinfinity):

geometrically this mean line y-2=x intersects the parabola y=(x+2)^2-6 in two points

OpenStudy (xapproachesinfinity):

no problem :)

OpenStudy (xapproachesinfinity):

let me show you what is happening Geometrically

OpenStudy (xapproachesinfinity):

http://prntscr.com/9ptxlx

OpenStudy (xapproachesinfinity):



OpenStudy (supernova_sonntag):

Okay I see :) Thanks again!

OpenStudy (xapproachesinfinity):

it's also important to see other interpretations